# Approximating Sparse Matrices and their Functions using Matrix-vector products

### v2

@@ 2.1.4 Adaptive methods
There are adaptive methods that reduce the number of matrix-vector products required for sparse recovery, for example [1, 43, 44, 50] by making adaptive measurements based on previous measurements. However, in our context we are concerned with recovering the whole matrix and it is difficult to make adaptive measurements that account for every row/column of $\bm{B}$ when we only have access to $\bm{B}$ using matrix-vector products. A naive strategy is to make adaptive measurements for each row/column, but this requires at least $\mathcal{O}(n)$ matrix-vector products, which is computationally infeasible and futile as we can recover $\bm{B}$ using $n$ measurements using the $n\times n$ identity matrix.
@@ 2.1.5 Shortcomings
There are some limitations with SpaMRAM. First, the entries of $f(\bm{A})$ are only negligible when $d(i,j)\gg 1$. Indeed, there are matrices for which the sparsity pattern implies $d(i,j)=\mathcal{O}(1)$ for all or most $i,j$. An example is the arrow matrix which has nonzero entries along the first row, first column and the main diagonal. The sparsity pattern of the arrow matrix makes $d(i,j)\leq 2$, which can make SpaMRAM behave poorly as the a priori estimate from (1) tells us that $f(\bm{A})$ may have no entries that are significantly larger than others. A similar issue has been discussed in [20], where the authors assume that the matrix we want to approximate is $d$-distributed, that is, the matrix has at most $d$ nonzero entries along any row or column. We will use a similar notion, which will be described in Subsection 2.2 to analyze SpaMRAM.
@@
We now define a notion similar to $d$-distributed sparse matrices in [20] to analyze SpaMRAM. Let $d\in\mathbb{Z}^{+}$ be a distance parameter and define
$k=\max_{1\leq i\leq n}\left\lvert\left\{j:d(i,j)\leq d\right\}\right\rvert.$
When $\bm{B}=f(\bm{A})$, $k$ measures the maximum number of vertices that are at most distance $d$ away from any given vertex. $k$ also measures the combined sparsity of $\bm{I}_{n},\bm{A},...,\bm{A}^{d}$ because the cardinality of the union of the support locations in a fixed row in each of $\bm{I}_{n},\bm{A},...,\bm{A}^{d}$ will be at most $k$. If $k=\mathcal{O}(1)$ for $d$ large enough such that $f$ can be approximated by a polynomial of degree $d-1$ on $\bm{A}$’s spectrum, then each row of $f(\bm{A})$ would only have $\mathcal{O}(1)$ dominating entries. Using the definition of $k$, we now analyze SpaMRAM in Theorem 1.
@@
for constants $C_{B}>0$, $01>\lambda_{B}>$ 0 and $d(i,j)\in\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is some bivariate function that captures the magnitude of the entries of $\bm{B}$. Then the output of SpaMRAM, $\widehat{\bm{B}}$, satisfies
$\left\lVert\widehat{\bm{B}}-\bm{B}\right\rVert_{2}\leq 9C_{B}\left(n+\frac{n^{3/2}}{\sqrt{k}}\right)\lambda_{B}^{d+1}$ (7)
@@
$k=\max_{i}\left\lvert\left\{j:d(i,j)\leq d\right\}\right\rvert$
@@
$\left\lVert E(i,:)\right\rVert_{2}\leq 9\left(\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{2}+\frac{1}{\sqrt{k}}\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{1}\right)$
where $\bm{b}_{i}$ is the $i$th row of $\bm{B}$ and $\bm{b}_{i}^{k}$ is the best $k$-term approximation to $\bm{b}_{i}$. Since $\bm{b}_{i}^{k}$ is the best $k$-term approximation to $\bm{b}_{i}$, we have
$\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{\max}\leq C_{B}\lambda_{B}^{d+1}$
@@
$\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{2}\leq\sqrt{n-k}C_{B}\lambda_{B}^{d+1},\qquad\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{1}\leq(n-k)C_{B}\lambda_{B}^{d+1}.$
@@
$\leq 9\sqrt{n}\max_{1\leq i\leq n}\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{2}+\frac{1}{\sqrt{k}}\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{1}$
$\leq 9\sqrt{n}\left(C_{B}\sqrt{n-k}\lambda_{B}^{d+1}+C_{B}\frac{n-k}{\sqrt{k}}\lambda_{B}^{d+1}\right)$
$<9C_{B}\left(n+\frac{n^{3/2}}{\sqrt{k}}\right)\lambda_{B}^{d+1}.$
@@
    $\left\lVert\bm{E}\right\rVert_{F}$ $\leq\sqrt{\sum_{1\leq i\leq n}\left\lVert E(i,:)\right\rVert_{2}^{2}}\leq\sqrt{n}\max_{1\leq i\leq n}\left\lVert E(i,:)\right\rVert_{2}<9C_{B}\left(n+\frac{n^{3/2}}{\sqrt{k}}\right)\lambda_{B}^{d+1}.$
@@
    The bound in (7) can be improved if for each $i$, the nonzero entries of $\bm{b}_{i}-\bm{b}_{i}^{k}$ decay exponentially. This implies
    $\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{1}\lesssim\lambda_{B}^{d+1},\qquad\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k}\right\rVert_{2}\lesssim\lambda_{B}^{d+1}$
@@
    $\left\lVert\bm{E}\right\rVert_{2}\lesssim\sqrt{n}\left(1+\frac{1}{\sqrt{k}}\right)\lambda_{B}^{d+1}.$
@@
    In general, the constants $C_{B},\lambda_{B}$ and the bivariate function $d(i,j)$ are either unknown, implicit or too expensive to compute in applications. For matrix functions, Theorem 1 relies on the decay bound by Benzi and Razouk [8], which contains constants that are implicit as they depend on the function $f$ and the spectral property of the matrix $\bm{A}$. Therefore, without a result that includes explicit constants, Theorem 1 cannot serve as a priori estimate for Algorithm 1. However, the analysis in Theorem 1 can still serve as a heuristic for when Algorithm 1 will be useful.
@@ 3 Approximation for functions of banded matrices
Suppose that $f$ is a function that is analytic in the region containing the spectrum of $\bm{A}$. We then have that the entries of $f(\bm{A})$ decay exponentially away from the main diagonal. More specifically, we have that there exists constants $C_{1},C_{2}>0$ and $10<\lambda_{1},\lambda_{2}<$ 1 dependent on $\bm{A}$ and $f$ such that for $i<j$,
@@
The proof of this result is based on polynomial approximation of $f$, together with the simple fact that powers of $\bm{A}$ are banded if $\bm{A}$ is. The exponents satisfy $\log\lambda_{1}^{-1}\propto k_{1}^{-1}$ and $\log\lambda_{2}^{-1}\propto k_{2}^{-1}$. For more details see [8, 9, 54]. Unlike the general sparse case, here the location of the large entries is known, i.e., near the main diagonal.
We will exploit the exponential off-diagonal decay given above in (8) and (9) to efficiently recover functions of banded matrices using matrix-vector products. We first design a way to recover banded matrices using matrix-vector products similarly to how banded Jacobian matrices were recovered in [16, 19] and then use the strategy used for the banded matrices along with the exponential off-diagonal decay to approximate functions of banded matrices and matrices with rapid off-diagonal decay using matrix-vector products. The idea of using the same strategy for matrices with rapid off-diagonal decay was suggested in [2] and was later applied to functions of banded matrices in [31].
@@ 3.1 Strategy for recovering banded matrices
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a banded matrix with upper bandwidth $k_{1}$ and lower bandwidth $k_{2}$. $\bm{A}$ has at most $1+k_{1}+k_{2}$ nonzero entries in each row and column, which lie consecutively near the main diagonal. Therefore if we recover the $1+k_{1}+k_{2}$ consecutive nonzero entries near the main diagonal in each row or column then we can recover $\bm{A}$. This motivates us to use a simple set of $s$ vectors, which to the authors’ knowledge, originates from [19] in the context of recovering matrices using matrix-vector products. The $s$ vectors are
@@
where $s<n$ is a positive integer and $\bm{I}_{s}(:,1:\{n/s\}\cdot s)$ is the first few columns of $\bm{I}_{s}$ to ensure $\bm{I}_{n}^{(s)}$ has exactly $n$ rows. This set of $s$ vectors capture $s$ consecutive nonzero entries of $\bm{A}$ in each row when right-multiplied to $\bm{A}$ if the other $(n-s)$ entries are all zeros. By choosing $s=1+k_{1}+k_{2}$ we are able to recover every nonzero entry of $\bm{A}$. We illustrate this idea with an example.
@@
We observe in the example that every nonzero entries of $\bm{A}_{6}$ in each row has been copied exactly once up to reordering onto the same row of $\bm{A}_{6}\bm{I}_{6}^{(4)}$. Since we know $\bm{A}_{6}$ has upper bandwidth $2$ and lower bandwidth $1$, we can reorganize the entries of $\bm{A}_{6}\bm{I}_{6}^{(4)}$ to recover $\bm{A}_{6}$ exactly. This idea can be generalized as follows [16, 19]. [^8] The work in [16, 19] is based on partitioning/coloring of the adjacency graph of a banded matrix. In this work, we show the vectors more explicitly.
Theorem 3.
@@
 [^9] The proof of Theorem 3 can be simplified by arguing that we need at most $1+k_{1}+k_{2}$ colors to color the columns of $A$ such that all the columns of the same color have disjoint support. Now, probing all the columns of the same color simultaneously yields the result. In this work, we provide a lengthier proof of Theorem 3 to make certain observations that lay the groundwork for our strategy to recover functions of banded matrices in Section 3.2.
@@ 3.2 Strategy for recovering functions of banded matrices
To accurately approximate $f(\bm{A})$ using matrix-vector products, we want to find a set of vectors when multiplied with $f(\bm{A})$, approximate the large entries of $f(\bm{A})$. We use the same set of vectors as for the banded case, namely $\bm{I}_{n}^{(s)}$. We assume that $s$ is an odd integer $s=2s_{0}+1$ throughout this section. We will see that this set of simple vectors can approximate functions of banded matrices. This is the same as the observation made in [2] and the work done later in [31], which are based on graph coloring of the adjacency graph of a matrix with off-diagonal decay.
@@
for some constant $C>0$ and $10<\lambda<$ 1. The results in this section can easily be extended to the more general case, that is, the class of matrices with entries satisfying (8) and (9), i.e., the decay rate along the two sides of the diagonal are different. See Subsection 3.2.4 for a further discussion.
The key idea is to revisit (11) in the proof of Theorem 3. For $1\leq i\leq n$ and $1\leq j\leq s$, recall that $t_{ij}$ is defined by
@@
$\left[\bm{B}\bm{I}_{n}^{(s)}\right]_{ij}=\underbrace{B_{i,j+st_{ij}}}_{\text{Dominant term}}+\underbrace{\sum_{\begin{subarray}{c}1\leq j+st\leq n\\ t\neq t_{ij}\in\mathbb{Z}\end{subarray}}B_{i,j+st}}_{\text{Sum of exponentially decaying terms}},$
@@
Figure 1 illustrates the error analysis given above for $\bm{B}=f(\bm{A})$ where $f$ is the exponential function $e^{x}$ and $\bm{A}\in\mathbb{R}^{1000\times 1000}$ is a symmetric $2$-banded matrix with entries drawn from i.i.d. $\mathcal{N}(0,n^{-1})$. The matrix-vector products $f(\bm{A})\bm{I}_{n}^{(s)}$ were evaluated using the polynomial Krylov method with $4,6$ and $8$ iterations (corresponding to polynomial approximation of degrees $3,5,7$ respectively), which give approximations to $f(\bm{A})\bm{I}_{n}^{(s)}$ of varying accuracy. In Figure 1a, we see the exponential decay in the entries of $f(\bm{A})$ and in Figure 1b, we see that (15) captures the decay rate until the error is dominated by the error from computing matrix-vector products using the Krylov method. The stagnation of the $\max\limits_{i=1,2,...,n}\left|[f(\bm{A})]_{i,i\pm(s-1)/2}\right|$ curve is caused by machine precision.
@@
The complexity of BaMRAM is $s$ matrix-vector products with $\bm{B}$. In the case $\bm{B}=f(\bm{A})$ for a $k$-banded matrix $\bm{A}$, we can evaluate $f(\bm{A})\bm{I}_{n}^{(s)}$ using polynomial or rational Krylov depending on $f$ and and the spectrum of $\bm{A}$. Assuming that the polynomial Krylov method converges (to sufficient accuracy) in $\mathcal{O}(1)$ iterations and $s=2ck+1$ for some constant $c$, the complexity of BaMRAM is $\mathcal{O}(snk)=\mathcal{O}(nk^{2})$. If the rational Krylov method is used, the complexity of BaMRAM is $\mathcal{O}(nk^{3})$.

### v3

@@ 2.1.4 Adaptive methods
There are adaptive methods that reduce the number of matrix-vector products required for sparse recovery, for example adaptive1; adaptive3; KrahmerWard2014; adaptive2 by making adaptive measurements based on previous measurements. However, in our context we are concerned with recovering the whole matrix and it is difficult to make adaptive measurements that account for every row/column of $\bm{B}$ when we only have access to $\bm{B}$ using matrix-vector products. A naive strategy is to make adaptive measurements for each row/column, but this requires at least $\mathcal{O}(n)$ matrix-vector products, which is computationally infeasible and futile as we can recover $\bm{B}$ using $n$ measurements using the $n\times n$ identity matrix.
@@ 2.1.5 Shortcomings
There are some limitations with SpaMRAM. First, the entries of $f(\bm{A})$ are only negligible when $d(i,j)\gg 1$. Indeed, there are matrices for which the sparsity pattern implies $d(i,j)=\mathcal{O}(1)$ for all or most $i,j$. An example is the arrow matrix which has nonzero entries along the first row, first column and the main diagonal. The sparsity pattern of the arrow matrix makes $d(i,j)\leq 2$, which can make SpaMRAM behave poorly as the a priori estimate from (1) tells us that $f(\bm{A})$ may have no entries that are significantly larger than others. A similar issue has been discussed in Dasarathyetal2015, where the authors assume that the matrix we want to approximate is $d$-distributed, that is, the matrix has at most $d$ nonzero entries along any row or column. We will use a similar notion, which will be described in Subsection 2.2 to analyze SpaMRAM.
@@
We now define a notion similar to $d$-distributed sparse matrices in Dasarathyetal2015 to analyze SpaMRAM. Let $d\in\mathbb{Z}^{+}$ be a distance parameter and define
$k_{d}=\max_{1\leq i\leq n}\left\lvert\left\{j:d(i,j)\leq d\right\}\right\rvert$.
When $\bm{B}=f(\bm{A})$, $k_{d}$ measures the maximum number of vertices that are at most distance $d$ away from any given vertex. $k_{d}$ also measures the combined sparsity of $\bm{I}_{n},\bm{A},...,\bm{A}^{d}$ because the cardinality of the union of the support locations in a fixed row in each of $\bm{I}_{n},\bm{A},...,\bm{A}^{d}$ will be at most $k_{d}$. If $k_{d}=\mathcal{O}(1)$ for $d$ large enough such that $f$ can be approximated by a polynomial of degree d1d-1 on $\bm{A}$’s spectrum, then each row of $f(\bm{A})$ would only have $\mathcal{O}(1)$ dominating entries. Using the definition of $k_{d}$, we now analyze SpaMRAM in Theorem 1.
@@
for constants $C_{B}>0$, $1>\lambda_{B}>0$ and $d(i,j)\in\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is some bivariate function that captures the magnitude of the entries of $\bm{B}$. Then the output of SpaMRAM, $\widehat{\bm{B}}$, satisfies
$\left\lVert\widehat{\bm{B}}-\bm{B}\right\rVert_{2}\leq 9C_{B}\left(n+\frac{n^{3/2}}{\sqrt{k_{d}}}\right)\lambda_{B}^{d+1}$ (7)
@@
$k_{d}=\max_{i}\left\lvert\left\{j:d(i,j)\leq d\right\}\right\rvert$
@@
$\left\lVert E(i,:)\right\rVert_{2}\leq 9\left(\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{2}+\frac{1}{\sqrt{k_{d}}}\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{1}\right)$
where $\bm{b}_{i}$ is the $i$th row of $\bm{B}$ and $\bm{b}_{i}^{k_{d}}$ is the best $k_{d}$-term approximation to $\bm{b}_{i}$. Since $\bm{b}_{i}^{k_{d}}$ is the best $k_{d}$-term approximation to $\bm{b}_{i}$, we have
$\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{\max}\leq C_{B}\lambda_{B}^{d+1}$
@@
$\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{2}\leq\sqrt{n-k_{d}}C_{B}\lambda_{B}^{d+1},\qquad\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{1}\leq(n-k_{d})C_{B}\lambda_{B}^{d+1}.$
@@
$\leq 9\sqrt{n}\max_{1\leq i\leq n}\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{2}+\frac{1}{\sqrt{k_{d}}}\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{1}$
$\leq 9\sqrt{n}\left(C_{B}\sqrt{n-k_{d}}\lambda_{B}^{d+1}+C_{B}\frac{n-k_{d}}{\sqrt{k_{d}}}\lambda_{B}^{d+1}\right)$
$<9C_{B}\left(n+\frac{n^{3/2}}{\sqrt{k_{d}}}\right)\lambda_{B}^{d+1}.$
@@
    $\left\lVert\bm{E}\right\rVert_{F}$ $\leq\sqrt{\sum_{1\leq i\leq n}\left\lVert E(i,:)\right\rVert_{2}^{2}}\leq\sqrt{n}\max_{1\leq i\leq n}\left\lVert E(i,:)\right\rVert_{2}<9C_{B}\left(n+\frac{n^{3/2}}{\sqrt{k_{d}}}\right)\lambda_{B}^{d+1}.$
@@
    The bound in (7) can be improved if for each $i$, the nonzero entries of $\bm{b}_{i}-\bm{b}_{i}^{k_{d}}$ decay exponentially. This implies
    $\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{1}\lesssim\lambda_{B}^{d+1},\qquad\left\lVert\bm{b}_{i}-\bm{b}_{i}^{k_{d}}\right\rVert_{2}\lesssim\lambda_{B}^{d+1}$
@@
    $\left\lVert\bm{E}\right\rVert_{2}\lesssim\sqrt{n}\left(1+\frac{1}{\sqrt{k_{d}}}\right)\lambda_{B}^{d+1}.$
@@
    In general, the constants $C_{B},\lambda_{B}$ and the bivariate function $d(i,j)$ are either unknown, implicit or too expensive to compute in applications. For matrix functions, Theorem 1 relies on the decay bound by Benzi and Razouk Benzi2007, which contains constants that are implicit as they depend on the function $f$ and the spectral property of the matrix $\bm{A}$. Therefore, without a result that includes explicit constants, Theorem 1 cannot serve as a priori estimate for Algorithm 1. However, the analysis in Theorem 1 can still serve as a heuristic for when Algorithm 1 will be useful.
@@ 3 Approximation for functions of banded matrices
Suppose that $f$ is a function that is analytic in the region containing the spectrum of $\bm{A}$. We then have that the entries of $f(\bm{A})$ decay exponentially away from the main diagonal. More specifically, we have that there exists constants $C_{1},C_{2}>0$ and $0<\lambda_{1},\lambda_{2}<1$ dependent on $\bm{A}$ and $f$ such that for $i<j$,
@@
The proof of this result is based on polynomial approximation of $f$, together with the simple fact that powers of $\bm{A}$ are banded if $\bm{A}$ is. The exponents satisfy $\log\lambda_{1}^{-1}\propto k_{1}^{-1}$ and $\log\lambda_{2}^{-1}\propto k_{2}^{-1}$. For more details see Benzi2007; BenziSimoncini2015; PozzaSimoncini2019. Unlike the general sparse case, here the location of the large entries is known, i.e., near the main diagonal.
We will exploit the exponential off-diagonal decay given above in (8) and (9) to efficiently recover functions of banded matrices using matrix-vector products. We first design a way to recover banded matrices using matrix-vector products similarly to how banded Jacobian matrices were recovered in ColemanMore1983; CPR1974 and then use the strategy used for the banded matrices along with the exponential off-diagonal decay to approximate functions of banded matrices and matrices with rapid off-diagonal decay using matrix-vector products. The idea of using the same strategy for matrices with rapid off-diagonal decay was suggested in BekasKokiopoulouSaad2007 and was later applied to functions of banded matrices in FrommerSchnimmelSchweitzer2021.
@@ 3.1 Strategy for recovering banded matrices
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a banded matrix with upper bandwidth $k_{1}$ and lower bandwidth $k_{2}$. $\bm{A}$ has at most $1+k_{1}+k_{2}$ nonzero entries in each row and column, which lie consecutively near the main diagonal. Therefore if we recover the $1+k_{1}+k_{2}$ consecutive nonzero entries near the main diagonal in each row or column then we can recover $\bm{A}$. This motivates us to use a simple set of $s$ vectors, which to the authors’ knowledge, originates from CPR1974 in the context of recovering matrices using matrix-vector products. The $s$ vectors are
@@
where $s<n$ is a positive integer and $\bm{I}_{s}(:,1:\{n/s\}\cdot s)$ is the first few columns of $\bm{I}_{s}$ to ensure $\bm{I}_{n}^{(s)}$ has exactly $n$ rows. This set of $s$ vectors capture $s$ consecutive nonzero entries of $\bm{A}$ in each row when right-multiplied to $\bm{A}$ if the other (ns)(n-s) entries are all zeros. By choosing $s=1+k_{1}+k_{2}$ we are able to recover every nonzero entry of $\bm{A}$. We illustrate this idea with an example.
@@
We observe in the example that every nonzero entry of $\bm{A}_{6}$ in each row has been copied exactly once up to reordering onto the same row of $\bm{A}_{6}\bm{I}_{6}^{(4)}$. Since we know $\bm{A}_{6}$ has upper bandwidth $2$ and lower bandwidth $1$, we can reorganize the entries of $\bm{A}_{6}\bm{I}_{6}^{(4)}$ to recover $\bm{A}_{6}$ exactly. This idea can be generalized as follows ColemanMore1983; CPR1974.999The work in ColemanMore1983; CPR1974 is based on partitioning/coloring of the adjacency graph of a banded matrix. In this work, we show the vectors more explicitly.
Proposition 3.
@@
 [^10] The proof of Proposition 3 can be simplified by arguing that we need at most $1+k_{1}+k_{2}$ colors to color the columns of $\bm{A}$ such that all the columns of the same color have disjoint support. Now, probing all the columns of the same color simultaneously yields the result. In this work, we provide a lengthier proof of Proposition 3 to make certain observations that lay the groundwork for our strategy to recover functions of banded matrices in Section 3.2.
@@ 3.2 Strategy for recovering functions of banded matrices
To accurately approximate $f(\bm{A})$ using matrix-vector products, we want to find a set of vectors when multiplied with $f(\bm{A})$, approximate the large entries of $f(\bm{A})$. We use the same set of vectors as for the banded case, namely $\bm{I}_{n}^{(s)}$. We assume that $s$ is an odd integer $s=2s_{0}+1$ throughout this section. We will see that this set of simple vectors can approximate functions of banded matrices. This is the same as the observation made in BekasKokiopoulouSaad2007 and the work done later in FrommerSchnimmelSchweitzer2021, which are based on graph coloring of the adjacency graph of a matrix with off-diagonal decay.
@@
for some constant $C>0$ and $0<\lambda<1$. The results in this section can easily be extended to the more general case, that is, the class of matrices with entries satisfying (8) and (9), i.e., the decay rate along the two sides of the diagonal are different. See Subsection 3.2.4 for a further discussion.
The key idea is to revisit (11) in the proof of Proposition 3. For $1\leq i\leq n$ and $1\leq j\leq s$, recall that $t_{ij}$ is defined by
@@
$\left[\bm{B}\bm{I}_{n}^{(s)}\right]_{ij}=\underbrace{B_{i,j+st_{ij}}}_{\text{Dominant term}}+\underbrace{\sum_{\begin{subarray}{c}1\leq j+st\leq n\\ t\neq t_{ij}\in\mathbb{Z}\end{subarray}}B_{i,j+st},}_{\text{Sum of exponentially decaying terms}}$
@@
Figure 1 illustrates the error analysis given above for $\bm{B}=f(\bm{A})$ where $f$ is the exponential function $e^{x}$ and $\bm{A}\in\mathbb{R}^{1000\times 1000}$ is a symmetric $2$-banded matrix with entries drawn i.i.d. from $\mathcal{N}(0,n^{-1})$. This matrix has real eigenvalues with $\lambda(A)\in[-0.1919,0.2011]$. The matrix-vector products $f(\bm{A})\bm{I}_{n}^{(s)}$ were evaluated using the polynomial Krylov method with $4,6$ and $8$ iterations (corresponding to polynomial approximation of degrees $3,5,7$ respectively), which give approximations to $f(\bm{A})\bm{I}_{n}^{(s)}$ of varying accuracy. In Figure 1a, we see the exponential decay in the entries of $f(\bm{A})$ and in Figure 1b, we see that (15) captures the decay rate until the error is dominated by the error from computing matrix-vector products using the Krylov method. The stagnation of the $\max\limits_{i=1,2,...,n}\left|[f(\bm{A})]_{i,i\pm(s-1)/2}\right|$ curve is caused by machine precision.
@@
The complexity of BaMRAM is $s$ matrix-vector products with $\bm{B}$. In the case $\bm{B}=f(\bm{A})$ for a $k$-banded matrix $\bm{A}$, we can evaluate $f(\bm{A})\bm{I}_{n}^{(s)}$ using Krylov method. Assuming that $f$ is analytic in a region containing the eigenvalues of $\bm{A}$, the polynomial Krylov method converges (to sufficient accuracy) in $\mathcal{O}(1)$ iterations and $s=2ck+1$ for some constant $c$, then the complexity of BaMRAM is $\mathcal{O}(snk)=\mathcal{O}(nk^{2})$.